# Routing in Quantum Networks with End-to-End Knowledge

### v1

@@ V-A Blocking Probability
Figure 7: Probability mass function (PMF) of path length with a fidelity threshold ($F_{th}$) = 0.53, and number of source and destination ($n_{sd}$) = 5.
Figure 8: Fairness vs. Fraction of high-quality nodes ($\xi$) in the transport network with a fidelity threshold ($F_{th}$) = 0.53 and number of source and destination ($n_{sd}$) = 5
Figure 9: Fidelity vs. Path-order ($\theta$) with a fidelity threshold ($F_{th}$) = 0.53, and number of source and destination ($n_{sd}$) = 12.
@@
Overall, the topology of the network has a visible effect on the performance of any approach. The blocking probability is relatively higher for regular topology as compared to random topology due to a shifted path length distribution in regular topology as seen in Figure 6. The regular structure of the network constrains the choice of routes which needs to deliver fidelity above the fidelity threshold. This results in a relatively smaller improvement in performance as compared to the case in random topology.
Figure 10: Fidelity vs. Path-order ($\theta$) with a fidelity threshold ($F_{th}$) = 0.53, and number of source and destination ($n_{sd}$) = 5.
@@ V-B Fairness
The majority of the approaches evaluated exhibit comparable levels of fairness and increases their fairness as $\xi$ increases, with the notable exception of the KSP approach, which demonstrates significantly lower fairness and decreases its fairness as $\xi$ increases. This means that, in general, these policies are able to exploit higher quality of repeaters in a fair manner. The reason for the drop of fairness experienced by KSP for high $\xi$ values is related to the corresponding increase of the blocking probabilities observed in Figure 4 i.e., KSP using unnecessarily long paths as shown in Figure 5 also for SD pairs that are topologically closer. In addition, those that are farther away (topologically) have thus a higher chance of being blocked, and this results in lower fairness values. Similarly, it is interesting to consider this result together with the performance in terms of blocking probability (Figure 4) even for slightly lower values of $\xi$ i.e., $\xi<0.8$. Although KSP has better BP in that range than SP, it does serve the non-blocked SD pairs fairly.
@@
The $kx_{0}$ approach not only maintains equitable treatment among users but also performs the best (among the considered approaches) in terms of BP (see Figure 4). Figure 9 ($n_{sd}=12$) and Figure 10 ($n_{sd}=5$) shows the fidelity distribution with respect to path order for different policies at $\xi=0.8$. For presentation purposes we only plot $kx_{0}$ path-selection from the ’grey box’ approaches leaving behind worse performing KSP and $kx_{1}$ path-selection as discussed above. It is to be noted that $\xi=0.8$ is chosen for analysis as the corresponding blocking probabilities are lower enough to be considered for realistic network settings. Since the policies are designed to deliver just over the threshold fidelity of 0.53, achieving higher fidelity than the fidelity threshold would be an over-utilisation of resources and can be considered as not fully optimised approach. The optimisation capability of $kx_{0}$ path-selection approach is visible from Figure 9 where it delivers a lower average fidelity (yet still over the fidelity threshold) for each path order with almost equal value as compared to other approaches. This shows that the approach uses low-quality nodes if the fidelity threshold allows for it which keeps the valuable resources to deliver the path for the upcoming path orders. The same is true for SP approach but the average fidelity values are more than the $kx_{0}$ path-selection approach which means that SP approach is not fully optimised. Contrary to this, KA has the worst case where it has unnecessarily higher average fidelity than the fidelity threshold for initial path orders and the average fidelity values gradually drops with path order. The same is true for $n_{sd}=5$ as well in Figure 10 but the effect is not clearly visible due to smaller number of source-destination pairs involved.
Figure 11: Fairness vs. Fraction of high-quality nodes ($\xi$) in the transport network with a fidelity threshold ($F_{th}$) = 0.53 and number of source and destination ($n_{sd}$) = 5
@@
Regular Topology: Figure 11 depicts the fairness in the regular topology. While all approaches exhibit comparable levels of fairness at $\xi=0.5$, the KA method demonstrates the least fairness as $\xi$ increases. This trend is attributable to the preferential consumption of high-quality nodes in the initial delivery paths, a phenomenon also evident from Figure 12, where the average fidelity decreases almost linearly with increasing path order, represented in orange. Contrary to this, it is to be noted that KA did not exhibit worst fairness in random topology as seen in Figure 8 due to relatively shorter path length as seen in Figure 5 than regular topology in Figure 7. The relatively longer path lengths, which arise from the inherent structure of the regular topology, enhance the effect of preferential consumption of high-quality nodes in regular topology. Although the blocking probability of the KA policy is closer to other policies as in Figure 6, it does not serve the paths fairly as in Figure 11.
@@
Figure 12: Fidelity vs. Path-order ($\theta$) with a fidelity threshold ($F_{th}$) = 0.53, and number of source and destination ($n_{sd}$) = 5.
Figure 13: Blocking probability (BP) vs. Fraction of high-quality nodes ($\xi$) in the transport network with a fidelity threshold ($F_{th}$) = 0.53 and number of source and destination ($n_{sd}$)
@@ V-C Number of source-destination pairs
Figure 4 ($n_{sd}=5$) and Figure 13 ($n_{sd}\in\{8,10,12\}$) collectively demonstrates the network’s performance under varying number of source-destination pairs for the random topology, displaying the blocking probability as a function of $\xi$. With increasing $n_{sd}$, the blocking probability escalates due to the limited availability of shared entangled pairs at quantum repeaters. Among the methods assessed, the $kx_{0}$ approach consistently outperforms all others across the different number of source-destination pairs. Notably, the $kx_{1}$ and KSP approaches, which perform well under lower number of source-destination pair, fail to maintain superiority over the SP approach as the $n_{sd}$ increases, exhibiting a decline in performance. This shows that the algorithm $kx_{i}$ is very sensitive to allowing paths even longer only 1 hop more than the shortest possible, and this is more and more visible as the $n_{sd}$ increases. At $n_{sd}=12$, the performance of SP, $kx_{0}$, and KA appears to converge, highlighting a uniform response to extreme network demands.
Figure 14: Blocking probability per edge (BP/E) vs. Fraction of high-quality nodes ($\xi$) in the transport network with a fidelity threshold ($F_{th}$) = 0.53, and number of source and destination ($n_{sd}$) = 5.
Figure 15: Blocking probability per edge (BP/E) vs. Fraction of high-quality nodes ($\xi$) in the transport network with a fidelity threshold ($F_{th}$) = 0.53, and number of source and destination ($n_{sd}$) = 5.
@@ V-D Average Degree
Figure 16: Blocking probability (BP) vs. Fraction of high-quality nodes ($\xi$) in the transport network with a fidelity threshold ($F_{th}$) = 0.53, and number of source and destination ($n_{sd}$) = 5.
@@ V-E Robustness
As described in Section III-A, the end-to-end fidelity data acquired through sampling is central to the $kx_{0}$ path-selection approach. In this aspect, we test the robustness of $kx_{0}$ approach by supplying inaccurate end-to-end fidelity data to the algorithm. We sample the end-to-end fidelity inaccuracy from a normal distribution given by:
@@
where $F_{1}$ is the correct end-to-end fidelity of the path; $F_{2}$ is the end-to-end fidelity of the path by replacing HQ repeater with a LQ one; $\lambda$ is a parameter modeling the end-to-end fidelity inaccuracy. Specifically, by varying $\lambda$, we modify the standard deviation of the noise added to the correct fidelity, thus controlling its magnitude.
Figure 16 presents the outcomes of a robustness assessment for the $kx_{0}$ approach for $\lambda\in\{1,4,8\}$ (thus, for decreasing magnitude of the noise). The blocking probability escalates as inaccuracies in the end-to-end fidelity data provided to the $kx_{0}$ approach increase or $\lambda$ decreases. Notably, even when $\lambda=1$, the performance of $kx_{0}$ remains superior to that of the SP approach. Therefore, our results suggest that the $kx_{0}$ approach exhibits robust performance under inaccurate fidelity estimations.
Figure 17: Blocking probability vs. Average degree ($\beta$) with a fidelity threshold ($F_{th}$) = 0.53, number of source and destination ($n_{sd}$) = 5, and having range of efficiency figures of quantum repeaters.
@@ V-F Range of Efficiency Figures
As a final aspect, we test the performance of the considered routing algorithms by relaxing the constraint on two node classes and assigning instead efficiency figures that are randomly drawn in the range given by eq 11, as detailed in Section IV. We set $a=10$ to attain practically useful performances of the routing policies. Setting a higher value of parameter $a$ would decrease the BP, even achieving BP $=0$ for higher $\beta$ values. Conversely, setting a lower value would increase the blocking probabilities above 0.5, which can be assumed to be unrealistic for practical purposes. The network configuration under consideration employs a random topology with $n_{sd}=5$. In both the $k$-shortest path and $kx$-path selection approaches, we test the performance for $k=10$ as well as $k=50$.
Figure 17 illustrates the correlation between the blocking probability and the Waxman parameter ($\beta$), with algorithm configurations i.e., $k=10$ (left) and $k=50$ (right) comprising nodes of varying efficiency. In comparison to the SP, all the other policies yielded superior results, with $kx_{0}$ achieving top performance, as already observed in the case of two-class quality repeaters.
With a lower $\beta$ value (i.e., lower number of links), some differences emerge, as explained in the following. The parameter $k$, that is the number of candidate shortest paths, has a minor influence on the performance, as illustrated in Figure 17. For a higher value i.e., $k=50$, the performance of $kx_{0}$ and $kx_{1}$ is marginally better compared to a lower value, i.e., $k=10$. In contrast, the performance of KSP slightly deteriorates when $k=50$. Although having additional shortest paths in the sample from which paths are selected does not seem beneficial for KSP , $kx_{0}$ and $kx_{1}$ benefit from the increased number of paths, even though the improvement is marginal. The reason is that KSP suffers from the same issue already discussed: the presence of additional paths leads to the selection of longer path lengths, which consumes valuable resources. Conversely, $kx_{0}$ efficiently utilises these additional paths without suffering from this drawback.

### v2

@@ V-A Blocking Probability
Figure 8: Probability mass function (PMF) of path length in a regular transport network with a fidelity threshold ($F_{th}$) of 0.53 and a source-destination count ($n_{sd}$) of 5.
Figure 9: Fairness as a function of the fraction of high-quality nodes ($\xi$) in a randomized transport network, with a fidelity threshold ($F_{th}$) of 0.53 and a source-destination count ($n_{sd}$) of 5.
Figure 10: Fidelity as a function of path-order ($\theta$) in a randomized transport network, with a fidelity threshold ($F_{th}$) of 0.53 and a source-destination count ($n_{sd}$) of 12.
@@
Overall, the network topology exerts a significant influence on the performance outcomes of each approach. Blocking probability is notably higher in the regular topology compared to the random topology, attributed to the shifted path length distribution in the regular topology (see Figure 6 & 8). This effect is evident in the comparative slopes of the blocking probability curves, with the regular topology in Figure 7 displaying a gentler slope than the random topology in Figure 5. The network’s regular structure limits the available route options because it offers little variation in served path lengths. This results in a relatively smaller improvement in performance as compared to the case in random topology.
Figure 11: Fidelity as a function of path-order ($\theta$) in a randomized transport network, with a fidelity threshold ($F_{th}$) of 0.53 and a source-destination count ($n_{sd}$) of 5.
Figure 12: Fairness as a function of the fraction of high-quality nodes ($\xi$) in a regular transport network, with a fidelity threshold ($F_{th}$) of 0.53 and a source-destination count ($n_{sd}$) of 5.
Figure 13: Fidelity as a function of path-order ($\theta$) in a regular transport network, with a fidelity threshold ($F_{th}$) of 0.53 and a source-destination count ($n_{sd}$) of 5.
@@ V-B Fairness
The majority of the approaches evaluated exhibit comparable levels of fairness and increases their fairness as $\xi$ increases, with the notable exception of the KSP approach, which demonstrates significantly lower fairness and decreases its fairness as $\xi$ increases. This means that, in general, these policies are able to exploit higher quality of repeaters in a fair manner. The reason for the drop of fairness experienced by KSP for high $\xi$ values is related to the corresponding increase of the blocking probabilities observed in Figure 5 i.e., KSP using unnecessarily long paths as shown in Figure 6 also for SD pairs that are topologically closer. In addition, those that are farther away (topologically) have thus a higher chance of being blocked, and this results in lower fairness values. Similarly, it is interesting to consider this result together with the performance in terms of blocking probability (Figure 5) even for slightly lower values of $\xi$ i.e., $\xi<0.8$. Although KSP has better BP in that range than SP, it does serve the non-blocked SD pairs fairly. The $kx_{0}$ approach not only maintains equitable treatment among users but also performs the best (among the considered approaches) in terms of BP (see Figure 5).
@@
To better illustrate user fairness from another perspective, the fidelity distribution for each path order ($\theta$)—representing the distribution of served fidelity for requests in sequence (i.e., 1 to $n_{sd}$)—is presented. Figure 10 ($n_{sd}=12$) and Figure 11 ($n_{sd}=5$) illustrate this fidelity distribution across path orders for various policies at $\xi=0.8$, with the average fidelity served for each path order also highlighted. For presentation purposes we only plot $kx_{0}$ path-selection from the ‘grey box’ approaches leaving behind worse performing KSP and $kx_{1}$ path-selection as discussed above. It is to be noted that $\xi=0.8$ is chosen for analysis as the corresponding blocking probabilities are lower enough to be considered for realistic network settings. Since the policies are designed to deliver just over the threshold fidelity of 0.53, achieving higher fidelity than the fidelity threshold would be an over-utilisation of resources and can be considered as not fully optimised approach. The optimisation capability of $kx_{0}$ path-selection approach is visible from Figure 10 where it delivers a lower average fidelity (yet still over the fidelity threshold) for each path order with almost equal value as compared to other approaches. This shows that the approach uses low-quality nodes if the fidelity threshold allows for it which keeps the valuable resources to deliver the path for the upcoming path orders. The same is true for SP approach but the average fidelity values are more than the $kx_{0}$ path-selection approach which means that SP approach is not fully optimised. Contrary to this, KA has the worst case where it has unnecessarily higher average fidelity than the fidelity threshold for initial path orders and the average fidelity values gradually drops with path order. The same is true for $n_{sd}=5$ as well in Figure 11 but the effect is not clearly visible due to smaller number of source-destination pairs involved.
@@
Regular Topology: Figure 12 illustrates the fairness within the regular topology, while Figure 13 shows the corresponding fidelity distribution for each path order ($\theta$) in the regular topology at $\xi=0.8$. The average fidelity served for each path order is highlighted, similar to the random topology case. While all approaches exhibit comparable levels of fairness at $\xi=0.5$, the KA method demonstrates the least fairness as $\xi$ increases. This trend is attributable to the preferential consumption of high-quality nodes in the initial delivery paths, a phenomenon also evident from Figure 13, where the average fidelity decreases almost linearly with increasing path order, represented in orange. Contrary to this, it is to be noted that KA did not exhibit worst fairness in random topology as seen in Figure 9 due to relatively shorter path length as seen in Figure 6 than regular topology in Figure 8. The relatively longer path lengths, which arise from the inherent structure of the regular topology, enhance the effect of preferential consumption of high-quality nodes in regular topology. Although the blocking probability of the KA policy is closer to other policies as in Figure 7, it does not serve the paths fairly as in Figure 12.
(a) $n_{sd}=8$
(b) $n_{sd}=10$
(c) $n_{sd}=12$
Figure 14: Blocking probability (BP) as a function of the fraction of high-quality nodes ($\xi$) in a randomized transport network with a fidelity threshold ($F_{th}$) of 0.53, shown for source-destination counts ($n_{sd}$) of: a) 8, b) 10, and c) 12.
@@ V-C Number of source-destination pairs
Figure 5 ($n_{sd}=5$) and Figure 14 ($n_{sd}\in\{8,10,12\}$) collectively demonstrates the network’s performance under varying number of source-destination pairs for the random topology, displaying the blocking probability as a function of $\xi$. With increasing $n_{sd}$, the blocking probability escalates due to the limited availability of shared entangled pairs at quantum repeaters. Among the methods assessed, the $kx_{0}$ approach consistently outperforms all others across the different number of source-destination pairs. Notably, the $kx_{1}$ and KSP approaches, which perform well under lower number of source-destination pair, fail to maintain superiority over the SP approach as the $n_{sd}$ increases, exhibiting a decline in performance. In the case of KSP, the BP begins to increase with $\xi$ when $\xi$ approaches values close to 1 (i.e., $\xi\in[0.9,1]$) for higher numbers of source-destination pairs ($n_{sd}\in\{8,10,12\}$). This counter-intuitive phenomenon arises from the KSP approach of selecting longer paths, which consumes valuable resources. This behaviour explained in Section V-A, is further intensified by the increased load resulting from a higher number of SD pairs operating within the network. This shows that the algorithm $kx_{i}$ is very sensitive to allowing paths even longer only 1 hop more than the shortest possible, and this is more and more visible as the $n_{sd}$ increases. At $n_{sd}=12$, the performance of SP, $kx_{0}$, and KA appears to converge, highlighting a uniform response to extreme network demands.
Table II consolidates all results obtained thus far, providing a comprehensive summary of the performance comparison across the evaluated approaches, organized by the transport network topologies. The table includes the performance metrics discussed in Sections V-A, V-B, and V-C, specifically blocking probability, fairness, and source-destination count.
Figure 15: Blocking probability per edge (BP/E) as a function of the fraction of high-quality nodes ($\xi$) in a randomized transport network, with a fidelity threshold ($F_{th}$) of 0.53 and a source-destination count ($n_{sd}$) of 5, for varying Waxman parameter values ($\beta$).
Figure 16: Blocking probability per edge (BP/E) as a function of the fraction of high-quality nodes ($\xi$) in a randomized transport network, with a fidelity threshold ($F_{th}$) of 0.53 and a source-destination count ($n_{sd}$) of 5, to model the performance of the $kx_{0}$ path-selection algorithm.
@@ V-D Average Degree
(a) $n_{sd}=5$
(b) $n_{sd}=8$
Figure 17: Robustness test: Blocking probability (BP) as a function of the fraction of high-quality nodes ($\xi$) in a randomized transport network, with a fidelity threshold ($F_{th}$) of 0.53 and a source-destination count ($n_{sd}$) of a) 5, b) 8. The robustness parameter ($\lambda$) models end-to-end fidelity inaccuracy, with $\lambda\in\{1,8,16\}$, corresponding to decreasing noise magnitude.
Figure 18: Robustness test: Blocking probability (BP) as a function of the fraction of high-quality nodes ($\xi$) in a regular transport network, with a fidelity threshold ($F_{th}$) of 0.53 and a source-destination count ($n_{sd}=5$). The robustness parameter ($\lambda$) models end-to-end fidelity inaccuracy, with $\lambda\in\{1,8,16\}$, corresponding to decreasing noise magnitude.
@@ V-E Robustness
As described in Section III-B, the end-to-end fidelity data acquired through sampling is central to the $kx_{0}$ path-selection approach. In this aspect, we test the robustness of $kx_{0}$ approach by supplying inaccurate end-to-end fidelity data to the algorithm. Specifically, we examine its performance on a random topology with $n_{sd}=\{5,8\}$ and a regular topology with $n_{sd}=5$. We sample the end-to-end fidelity inaccuracy from a normal distribution given by:
@@
where $F_{1}$ is the correct end-to-end fidelity of the path; $F_{2}$ is the end-to-end fidelity of the path by replacing a HQ repeater with a LQ one; $\lambda$ is a parameter modelling the end-to-end fidelity inaccuracy. Specifically, by varying $\lambda$, we modify the standard deviation of the noise added to the correct fidelity, thus controlling its magnitude. We classify these as follows: $\lambda=1$ (High inaccuracy), $\lambda=8$ (Moderate inaccuracy), and $\lambda=16$ (Low inaccuracy).
Figure 17a for $n_{sd}=5$ and  17b for $n_{sd}=8$ presents the outcomes of a robustness assessment for the $kx_{0}$ approach for $\lambda\in\{1,8,16\}$ in a random network. The blocking probability escalates as inaccuracies in the end-to-end fidelity data provided to the $kx_{0}$ approach increase or $\lambda$ decrease. For $n_{sd}=5$, even introduction of high inaccuracy, the performance of $kx_{0}$ remains superior to that of the SP approach. The $kx_{0}$ continue to better SP under demanding condition i.e., with more SD pairs ($n_{sd}=8$) in low and moderate inaccuracies cases. However, under more demanding conditions i.e., when fidelity inaccuracies are high along with a higher number of source-destination pairs ($n_{sd}=8$), the advantage of $kx_{0}$ over SP diminishes, as shown in Figure 17b. Similarly, Figure 18 shows the robustness results for the $kx_{0}$ approach in a regular network for $\lambda=\{1,8,16\}$. Consistent with the random topology results, $kx_{0}$ outperforms SP approach even at high fidelity inaccuracies, although the performance differences between accuracy levels decrease. This reduced performance variation is due to the regular network structure, which constrains path diversity between SD pairs as compared to the random network.
Overall, these results suggest that the $kx_{0}$ approach demonstrates robust performance with inaccurate fidelity estimates, except under highly demanding conditions.
(a) For $k=10$
(b) For $k=50$
Figure 19: Blocking probability vs. Average degree ($\beta$) with a fidelity threshold ($F_{th}$) = 0.53, number of source and destination ($n_{sd}$) = 5 in a random topology, and having range of efficiency figures of quantum repeaters. Comparison of shortest-path with grey-box algorithms having algorithm configuration parameter: a) $k=10$, b) $k=50$.
@@ V-F Range of Efficiency Figures
As a final aspect, we test the performance of the considered routing algorithms by relaxing the constraint on two node classes and assigning instead efficiency figures that are randomly drawn in the range given by eq 11, as detailed in Section IV. We set $a=10$ to attain practically useful performances of the routing policies. Setting a higher value of parameter $a$ would decrease the BP, even achieving BP $=0$ for higher $\beta$ values. Conversely, setting a lower value would increase the blocking probabilities above 0.5, which can be assumed to be unrealistic for practical purposes. The network configuration under consideration employs a random topology with $n_{sd}=5$. In both the $k$-shortest path (with $k=10$ as $k_{10}$) and $kx$-path selection (with $k=10,\ x=0$ as $k_{10}x_{0}$; $k=10,\ x=1$ as $k_{10}x_{1}$; $k=50,\ x=0$ as $k_{50}x_{0}$; $k=50,\ x=1$ as $k_{50}x_{1}$) approaches, we test the performance for $k=10$ as well as $k=50$.
Figure 19 illustrates the correlation between the blocking probability and the Waxman parameter ($\beta$), with algorithm configurations i.e., $k=10$ (Figure 19a) and $k=50$ (Figure 19b) comprising nodes of varying efficiency. In comparison to the SP, all the other policies yielded superior results, with $kx_{0}$ achieving top performance, as already observed in the case of two-class quality repeaters.
With a lower $\beta$ value (i.e., lower number of links), some differences emerge, as explained in the following. The parameter $k$, that is the number of candidate shortest paths, has a minor influence on the performance, as illustrated in Figure 19. For a higher value i.e., $k=50$, the performance of $kx_{0}$ and $kx_{1}$ is marginally better compared to a lower value, i.e., $k=10$. In contrast, the performance of KSP slightly deteriorates when $k=50$. Although having additional shortest paths in the sample from which paths are selected does not seem beneficial for KSP. $kx_{0}$ and $kx_{1}$ benefit from the increased number of paths, even though the improvement is marginal. The reason is that KSP suffers from the same issue already discussed: the presence of additional paths leads to the selection of longer path lengths, which consumes valuable resources. Conversely, $kx_{0}$ efficiently utilises these additional paths without suffering from this drawback.